OpenStudy (anonymous):

Make an (x,y) Table in which is there is a constant ratio of -1/4 between the x- and y- values. include at least four pairs of corresponding values. 1 is done. here 1. x: -8, -4 , 4 , 8 y: 2 , 1 , -1 , -2 2.make a prediction: will the graph of this relationship slope up or down to the right? 3. As x increases by 1 unit, by how much does y increase or decrease? this value is the slope of the line. 4. sketch a graph of this relationship. HELP ME PLS

OpenStudy (anonymous):

i would say it would go down, because as \(x\) gets larger, \(y\) gets smaller

OpenStudy (anonymous):

because the constant ratio is \(-\frac{1}{4}\), as \(x\) increases by 1 unit, \(y\) decreases by \(\frac{1}{4}\) unit

OpenStudy (anonymous):

well the ratio is \[-\frac{1}{4}\] sooo in that sense.. the slope is also the ratio....